The continuous signal f(t) = cos(πt/2) sampled at 1 second intervals starting from t = 0.


(a) Find the Laplace transform of the sampled signal f*(t)


1
Expert's answer
2021-07-14T10:10:35-0400

"f^*(s)=\int\limits_0^1\cos\frac{\pi t}{2}e^{-ts}dt=-\frac{1}{s}\int\limits_0^1\cos\frac{\pi t}{2}de^{-ts}="

"[-\frac{1}{s}\cos\frac{\pi t}{2}e^{-ts}]_0^1+\frac{1}{s}\int\limits_0^1e^{-ts}d\cos\frac{\pi t}{2}=" "\frac{1}{s}-\frac{\pi}{2s}\int\limits_0^1\sin\frac{\pi t}{2}e^{-ts}dt"


"\int\limits_0^1\sin\frac{\pi t}{2}e^{-ts}dt=-\frac{1}{s}\int\limits_0^1\sin\frac{\pi t}{2}de^{-ts}=""[-\frac{1}{s}\sin\frac{\pi t}{2}e^{-ts}]_0^1+\frac{1}{s}\int\limits_0^1e^{-ts}d\sin\frac{\pi t}{2}="

"=-\frac{1}{s}e^{-s}+\frac{\pi}{2s}\int\limits_0^1\cos\frac{\pi t}{2}e^{-ts}dt=-\frac{1}{s}e^{-s}+\frac{\pi}{2s}f^*(s)"


Hence, "f^*(s)=\frac{1}{s}-\frac{\pi}{2s}\left(-\frac{1}{s}e^{-s}+\frac{\pi}{2s}f^*(s)\right)"

"(1+\frac{\pi^2}{4s^2})f^*(s)=\frac{1}{s}+\frac{\pi}{2s^2}"


Therefore, "f^*(s)=\frac{\frac{1}{s}+\frac{\pi}{2s^2}}{1+\frac{\pi^2}{4s^2}}=\frac{2\pi+4s}{\pi^2+4s^2}"